OpenStudy (anonymous):

UESTION 1 c=0.05x^2-2x+50 The cost for producing stuffed animals is modeled by the function . Which of the following is true? The minimum cost occurs when 20 stuffed animals are produced. The minimum cost occurs when 30 stuffed animals are produced. The minimum cost is $20. The maximum cost is $20.

OpenStudy (anonymous):

anyone know

OpenStudy (anonymous):

0.05 times itself is 0.0025

OpenStudy (anonymous):

+50 is 50.0025

OpenStudy (anonymous):

is it a b c or d

OpenStudy (anonymous):

divided by 2 is 25 ish

OpenStudy (anonymous):

ok so c?

OpenStudy (welshfella):

you can use calculus to do this Have wou studied calculus?

OpenStudy (anonymous):

no im in algebra 2

OpenStudy (anonymous):

caculus next year

OpenStudy (anonymous):

yea i think ur correct

OpenStudy (welshfella):

oh right well yo can do its by either drawing a graph or converting the function to vertex form

OpenStudy (anonymous):

so b

OpenStudy (welshfella):

Can you complete the square?

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

right so proceed as follows 0.05x^2 - 2x + 50 = 0.05 ( x^2 - 40x + 1000) = 0.05[ (x - 20)^2 + 1000 - 400) thats the vertex form this has a minimum value because coefficient of x^2 is positive and at the minimum point x = 20

OpenStudy (welshfella):

oh well!!